OpenStudy (anonymous):

Beth is x years old today,and Jemaine is four times as old as Beth. In four years,Jermaine will be twice as old as Beth.How old is Jermaine today? 1)2 2)4 3)6 4)8 5)12

OpenStudy (unklerhaukus):

B=x J=4x 2.(B+4)=J+4

OpenStudy (anonymous):

okay

OpenStudy (unklerhaukus):

Beth is x years old today, and Jemaine is four times as old as Beth B = x , J = 4 B = 4 x

OpenStudy (unklerhaukus):

In four years, Jermaine will be twice as old as Beth +4 ; J+4 = 2 ( B+4 )